# Language Agents for Hypothesis-driven Clinical Decision Making with Reinforcement Learning

### v1

@@ 5 Experimental Set-up
5.1 Dataset and Pre-Processing
We evaluate our method on the MIMIC-CDM dataset Hager et al. (2024), a curated subset of MIMIC-IV Johnson et al. (2020) designed for modeling sequential clinical decision making. It contains 2,400 patients diagnosed with one of four abdominal conditions: appendicitis, cholecystitis, diverticulitis, or pancreatitis. This focused setting on four pathologies reflects real clinical workflows, where physicians perform differential diagnosis with a narrowed down space of possible diseases and request tests to distinguish between likely candidates. The dataset includes patient histories (symptoms, comorbidities, family histories) and physical exam notes for all patients. It also provides 5,959 textual imaging reports (CT, x-ray, ultrasound, MRI) and 143,191 lab results (blood, urine, microbiology), however, not every test result is reported for every patient. If multiple values for a specific test were recorded during the hospital stay of the patient only the first one was included in MIMIC-CDM to simulate an early diagnosis after hospital admission. Crucially, the dataset includes comprehensive mappings of test names across patients, an essential feature for modeling test requests reliably. Without this normalization, a model could not query the same test across different cases due to inconsistent naming in clinical documentation. To our knowledge, MIMIC-CDM is the only publicly available dataset that enables simulation of this setting.
For our use-case, we construct the set of available tests as: physical examination, all imaging modalities, and the most common laboratory panels. These panels are collections of individual tests that are usually ordered together. The initial patient history is shortened by processing them with a Mixtral-8x7B Jiang et al. (2024) model, prompted to summarize the most important aspects. Also the imaging reports are shortened, wherever a separation into sections was available, by only keeping the findings section of the report and removing the remaining sections. The original dataset does not have a data split, since it was intended for the evaluation of zero-shot models. We therefore split the data into a training set of 80%, and a validation and test set of 10% each.
5.2 Metrics
To evaluate model performance, we report class-wise accuracies along with their mean, as well as micro and macro F1-scores. As classification is performed by auto-regressive LLMs, some generations do not include a valid class. If this is the case, we fall back to the last classification of the hypothesis agent. If the method under evaluation does not have a hypothesis agent, or that hypothesis agent also predicts an invalid class, those predictions are defined as predictions of a separate fifth wrong class, however, we exclude their contribution to the macro F1-score.
Additionally, we compute the Expected Calibration Error (ECE) to assess confidence calibration of our hypothesis agent. ECE measures the discrepancy between predicted confidence scores and actual accuracy. A lower ECE indicates better model calibration, meaning the predicted probabilities align well with actual correctness frequencies.
@@ 5.2 Metrics
Method Append. Cholec. Divert. Pancr. Mean Micro Macro # Tests
OASST Hager et al. (2024)* 82.0 48.0 45.5 44.1 54.9 - - -
SFT-all 89.6 78.5 96.0 66.6 82.7 82.1 82.4 9.61 (all)
SM-DDPO Yu et al. (2023) 74.3 0.0 15.6 58.0 37.0 45.4 31.8 6.46
ReAct Yao et al. (2022) 57.3 67.7 76.0 38.9 60.0 57.9 46.0 2.62
LA-CDM (ZS) 67.7 56.9 72.0 44.4 60.3 60.4 56.8 2.25
LA-CDM 93.8 80.0 60.0 59.3 73.3 80.4 77.4 0.65
Table 1: Performance comparison of LA-CDM and baseline methods. We report class-wise accuracies and F1-scores. The avg. # tests is the mean number of tests that were requested by the model for all patients in the test set. *OASST is evaluated on a different test set in a different framework. SM-DDPO can only process tabular data. ZS = zero-shot.
5.3 Baselines
We compare our method to various zero-shot and trained baselines. In their work, Hager et al. (2024) introduce both the MIMIC-CDM dataset and evaluate how various "out-of-the-box" pre-trained models perform when tasked with clinical decision-making. We compare with OASST, the best performing model from this evaluation, however, a direct comparison is very difficult for multiple reasons. First, as a zero-shot method, the model was evaluated on the complete dataset, whereas we reserve some part of that dataset for training. Second, the clinical decision framework was constructed differently, as the OASST model is not provided with the possible diagnosis classes or available tests in its prompt. Additionally, as an approximate upper-bound for our method, we compare with a Llama-3-7B-Instruct model Dubey et al. (2024) identical to our LLM backbone, trained with supervised fine-tuning to predict the correct diagnosis. Instead of requesting tests it simply receives all the available patient information directly which is not a realistic diagnostic process. We refer to this method as SFT-all. We compare to three other adaptive test selection methods on MIMIC-CDM. SM-DDPO Yu et al. (2023) trains a MLP using reinforcement learning for clinical decision making, however, the method is only able to process tabular data and can only request laboratory values. ReAct Yao et al. (2022) is a zero-shot decision making method. Finally, we compare with an untrained zero-shot version of our method, LA-CDM (ZS), including hypothesis and decision agent.
@@ 6 Results and Discussion
HA DA Append. Cholec. Divert. Pancr. Mean Micro Macro Avg. # Tests
$\checkmark$ 88.5 72.3 48.0 44.4 63.3 70.0 73.6 2.20
$\checkmark$ $\checkmark$ 93.8 80.0 60.0 59.3 73.3 80.4 77.4 0.65
Table 2: Ablation study of the hypothesis-driven approach. HA = hypothesis agent, DA = decision agent.
6.1 Comparison with Baselines
@@
Our comparison with baselines is shown in Table 1. When comparing with OASST Hager et al. (2024), LA-CDM shows improvement in accuracy for each class resulting in almost 20 percentage points difference when comparing the mean of all class accuracies. The performance shows the largest improvement for pathologies that are less common compared with the majority class appendicitis. While the methods are not directly comparable as outlined in section 5.3, the high performance improvement clearly shows the advantage of training on the task of clinical decision-making and not relying on inherent capabilities of large pre-trained models. This is further supported by the evaluation results of our method as a zero-shot model. In comparison to our trained model, we see a great performance improvement achieved through training. Importantly, the zero-shot method also diagnoses less efficiently than the trained version, requiring on average almost 3.5x more tests to form a diagnosis. Also the individual performance of the hypothesis agent benefits from training. Through our hypothesis generation training, we improve the ability of the model to form correct hypotheses from 54.6% to 74.2%. We equally show an improvement of uncertainty-awareness through training the confidence calibration objective. The ECE decreases significantly from 0.226 to 0.150. We visualise the calibration curves and confidence distribution of the two models in Figure 3, where the trained model shows better calibration, especially at often predicted confidences. It also shows more distributed confidence predictions, compared to the untrained model, which tends to be very over-confident, mostly predicting a confidence of 80%.
The SFT-all model serves as a rough upper bound, as it leverages all available retrospective patient data, an unrealistic setup for real-time clinical decision-making. Therefore, it cannot be used for direct patient interactions. LA-CDM performs comparably, trailing by only two and five percentage points in micro and macro F1-score, respectively, while requiring 15x fewer diagnostic tests. This substantial reduction in test count highlights the efficiency of our approach.
Moreover, we observe evidence of patient-adaptive testing strategies aligning with best practices: for suspected cholecystitis, the model most frequently selects ultrasound (32% of cases), the gold-standard test Hirota et al. (2007); for appendicitis, it prioritizes CT scans (8% of cases), consistent with diagnostic guidelines Di Saverio et al. (2020). These results demonstrate that our method not only achieves high diagnostic accuracy but also optimizes resource usage in a clinically meaningful way. The reasoning traces generated by chain-of-thought prompting further enable the interpretation of the model’s testing pathways. We report qualitative examples of the model’s generations in Appendix C.
6.2 Ablation Study
@@
We evaluate the benefit of our hypothesis-driven approach in Table 2. When removing the hypothesis agent from our methodology, the decision agent has to learn to request tests and to diagnose without relying on the uncertainty-aware hypothesis generation capabilities of the hypothesis agent, explicitly trained for these objectives. The benefit of the hypothesis agent is demonstrated clearly by improvements in all metrics. Notably, including the hypothesis agent also significantly reduces the amount of tests requested by the model.
6.3 Limitations
@@
This work presents a first step towards training LLMs for clinical decision-making. As such there are some limitations that should be addressed in future work. First, the data we are training on is limited. It only contains four abdominal pathologies and a limited number of available diagnostic tests, the model is therefore only trained in clinical decision-making for these diseases. An extension to more diseases and more tests remains as future work. Secondly, the data we are training on is retrospective with different tests missing for different patients. Furthermore, the available tests are those tests that the clinicians involved in treating that patient performed. The model can only explore a limited spread of testing pathways. It can therefore only learn to become more efficient within the testing protocols performed by doctors. Simulation of unavailable test data could open up a pathway for modeling a more holistic clinical decision-making environment.
@@ A.1 Hypothesis Agent Prompt
You are an uncertainty-aware and helpful medical expert. Your task is to create a hypothesis for a diagnosis of a patient admitted to the emergency department complaining about abdominal symptoms. You can hypothesise one of the following four conditions:
@@
Additionally, indicate your confidence using a confidence score between 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10. A value close to 0 means you think there is a high probability that the answer is wrong. The closer the value is to 10, the higher you think is the probability that the answer is correct. Be very fine-grained with your confidence estimation and make use of the whole number range.
Always use the following format for your answer:
@@
Confidence: Here, you put your confidence in this hypothesis on a scale of 0 to 10.
Here are the four master rules you need to follow:
1. Always provide a hypothesis, even if you are uncertain.
2. Always provide a confidence with full integers from 0 to 10.
3. Only give one of the four possible conditions as hypothesis and use their exact naming.
4. Do not answer with anything other than the hypothesis and confidence, do not give explanations, and follow the exact format given above.
@@
Here is some information on the patient:
{observed patient state}
@@ A.2 Decision Agent Prompt
To aid you in your diagnosis, you can request the results for the following tests for the patient:
@@
Thought: Here, you should think about what to do next. Be very brief.
Action: Here, you can either write "Diagnosis" if you are confident enough to provide a diagnosis or "Test" if you want to request a single test.
Action Input: If you chose diagnosis as your action, you should put the correct condition of the four provided above here. You MUST put a condition here. If you chose test, provide the test name from the list mentioned above. You can only request one single test per action. Use the exact naming of the test as given above. Do not add any specification, abbreviation or explanation to the test name. If you do this there will be detrimental consequences to the patient.
$<$submit$>$
If you requested a test, the user will provide you with the test results if available. Do not hallucinate the test results yourself. Don’t write any more text after that. If the test is not available, do not request the same test again.
Here are the four master rules:
@@
2. Only ask for one of the tests given in the list above and do not ask for any test twice.
@@
5. Keep asking for tests until you are sufficiently confident but always provide a valid diagnosis in the end.
@@
Here are some examples:
1)
Thought: Given the symptoms, I should order a Physical Examination first.
Action: Test
Action Input: Physical Examination
$<$submit$>$
2)
Thought: The CT shows clear inflammation of the appendix.
Action: Diagnosis
Action Input: appendicitis
$<$submit$>$
@@
{observed patient state}
The current hypothesis is {hypothesis} with confidence {hypothesis confidence}.
Appendix B Experiments with different LLM backbones
To show model family variability, we show performance comparisons of Qwen 2 7B instruct (https://huggingface.co/Qwen/Qwen2-7B-Instruct) and Llama 3.1 8B instruct (https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct). The Qwen 2 model performs similarly to its Llama 3.1 counterpart, without changing any hyperparameters, highlighting the robustness of our approach. When evaluating smaller model scales, e.g., Qwen 2 1.5B instruct, we found a strong deterioration in performance and fomat-keeping difficulties, both zero-shot (ZS) and after training, showing the task’s difficulty. We report the results for the Qwen 2 8B instruct model in the Table 3.

### v2

@@ 5 Experimental Set-up
Dataset and Pre-Processing
We evaluate our method on the MIMIC-CDM dataset (Hager et al., 2024), a curated subset of MIMIC-IV (Johnson et al., 2020) designed for modeling sequential clinical decision making. It contains 2,400 patients diagnosed with one of four abdominal conditions: appendicitis, cholecystitis, diverticulitis, or pancreatitis. This focused setting on four pathologies reflects real clinical workflows, where physicians perform differential diagnosis with a narrowed down space of possible diseases and request tests to distinguish between likely candidates. The dataset includes patient histories (symptoms, comorbidities, family histories) and physical exam notes for all patients. It also provides 5,959 textual imaging reports (CT, x-ray, ultrasound, MRI) and 143,191 lab results (blood, urine, microbiology), however, not every test result is reported for every patient. If multiple values for a specific test were recorded during the hospital stay of the patient only the first one was included in MIMIC-CDM to simulate an early diagnosis after hospital admission. Crucially, the dataset includes comprehensive mappings of test names across patients, an essential feature for modeling test requests reliably. Without this normalization, a model could not query the same test across different cases due to inconsistent naming in clinical documentation. To our knowledge, MIMIC-CDM is the only publicly available dataset that enables simulation of this setting. For our use-case, we construct the set of available tests as: physical examination, all imaging modalities, and the most common laboratory panels. These panels are collections of individual tests that are usually ordered together.
Metrics
To evaluate model performance, we report class-wise accuracies along with their mean, as well as micro and macro F1-scores. Additionally, we compute the Expected Calibration Error (ECE) to assess confidence calibration of our hypothesis agent. ECE measures the discrepancy between predicted confidence scores and actual accuracy. A lower ECE indicates better model calibration, meaning the predicted probabilities align well with actual correctness frequencies.
Table 1: Performance comparison of LA-CDM and baseline methods. We report class-wise accuracies and F1-scores. The avg. test cost is the mean cost of the tests for each patients in the test set. *OASST results are taken from Hager et al. (2024). It is evaluated on a different test set in a different framework. SM-DDPO can only process tabular data. ZS = zero-shot
@@
Method Append. Cholec. Divert. Pancr. Mean Micro Macro Test Cost
OASST* 82.0 48.0 45.5 44.1 54.9 - - -
SFT-all 98.4 89.8 95.8 87.5 92.8 93.6 92.9 \$3792.79
SM-DDPO 74.3 0.0 15.6 58.0 37.0 45.4 31.8 -
ReAct 90.2 79.7 66.7 62.9 74.9 79.1 74.8 \$1480.32
LA-CDM (ZS) 73.5 55.1 72.0 57.5 64.5 65.3 64.5 \$1521.73
LA-CDM 93.1 83.6 75.0 73.5 81.3 84.1 81.3 \$1295.61
Baselines
We compare our method to various zero-shot and trained baselines. In their work, Hager et al. (2024) introduce both the MIMIC-CDM dataset and evaluate how various ”out-of-the-box” pre-trained models perform when tasked with clinical decision-making. We compare with OASST, the best performing model from this evaluation, however, a direct comparison is very difficult for multiple reasons. First, as a zero-shot method, the model was evaluated on the complete dataset, whereas we reserve some part of that dataset for training. Second, the clinical decision framework was constructed differently, as the OASST model is not provided with the possible diagnosis classes or available tests in its prompt. Additionally, as an approximate upper-bound for our method, we compare with a Qwen-2.5-7B-Instruct model (Yang et al., 2024) identical to our LLM backbone, trained with supervised fine-tuning to predict the correct diagnosis. Instead of requesting tests it simply receives all the available patient information directly which is not a realistic diagnostic process. We refer to this method as SFT-all. We compare to three other adaptive test selection methods on MIMIC-CDM. SM-DDPO (Yu et al., 2023) trains a MLP using reinforcement learning for clinical decision making, however, the method is only able to process tabular data and can only request laboratory values. ReAct (Yao et al., 2022) is a zero-shot decision making method and LA-CDM (ZS) is the untrained version of our method.
@@ 6 Results and Discussion
Table 2: Ablation of the inclusion of the test cost reward function.
Accuracy F1-Score
Test Cost Append. Cholec. Divert. Pancr. Mean Micro Macro Avg. Test Cost
93.5 81.4 82.1 72.3 82.3 84.3 82.4 \$1427.85
$\checkmark$ 93.1 83.6 75.0 73.5 81.3 84.1 81.3 \$1295.61
Table 3: Ablation of the hypothesis-driven approach. Experiments without test cost. HA = hypothesis agent, DA = decision agent.
@@
HA DA Append. Cholec. Divert. Pancr. Mean Micro Macro Avg. Test Cost
$\checkmark$ 93.0 79.7 79.2 62.2 78.5 81.7 78.6 \$1410.01
$\checkmark$ $\checkmark$ 93.5 81.4 82.1 72.3 82.3 84.3 82.4 \$1427.85
Comparison with Baselines
@@
Our comparison with baselines is shown in Table 1. When comparing with OASST (Hager et al., 2024), LA-CDM shows improvement in accuracy for each class resulting in almost 30 percentage points difference when comparing the mean of all class accuracies. The performance shows the largest improvement for pathologies that are less common compared with the majority class appendicitis. While the methods are not directly comparable as outlined in section 5, the high performance improvement clearly shows the advantage of training on the task of clinical decision-making and not relying on inherent capabilities of large pre-trained models. This is further supported by the evaluation results of ReAct and of LA-CDM as a zero-shot model. In comparison to our trained model, we see a great performance improvement achieved through training. Importantly, the untrained methods also diagnoses less efficiently than the trained version, acquiring, in the case of ReAct, almost \$200 more test cost to form a diagnosis. Also the individual performance of the hypothesis agent benefits from training LA-CDM. Through our hypothesis generation training, we improve the ability of the model to form correct hypotheses from 75.7% to 81.9%. We equally show an improvement of uncertainty-awareness through training the confidence calibration objective. The ECE decreases from 0.069 to 0.037. We visualise the calibration curves and confidence distribution of the two models in Figure 3, where the trained model shows better calibration, especially at often predicted confidences. When comparing with SM-DDPO, the other trained baseline, we can see the benefit of being able to process textual input. SM-DDPO almost entirely fails as it is limited to tabular data which is not sufficient in many diagnostic tasks like the one at hand, where imaging results are paramount. The SFT-all model serves as a rough upper bound, as it leverages all available retrospective patient data, an unrealistic setup for real-time clinical decision-making. Therefore, it cannot be used for direct patient interactions. Naturally, the performance is very strong, however, as all available tests are used, the diagnostic cost is more than tripled. The substantial reduction in test cost highlights the efficiency of our approach. Moreover, we observe evidence of patient-adaptive testing strategies aligning with best practices: for suspected cholecystitis, the model most frequently selects ultrasound (64.9% of cases), the gold-standard test (Hirota et al., 2007); for appendicitis, it prioritizes CT scans (85.1% of cases), consistent with diagnostic guidelines (Di Saverio et al., 2020). These results demonstrate that our method not only achieves high diagnostic accuracy but also optimizes resource usage in a clinically meaningful way. The reasoning traces generated by chain-of-thought prompting further enable the interpretation of the model’s testing pathways. We report qualitative examples of the model’s generations in Appendix B.
Ablation Study
@@
In Table 2, we observe that while the inclusion of the test cost reward function, performs similarly to training without test cost, the average test cost per patient is reduced significantly. This shows how the model learns to become more efficient in its testing pathways while achieving similar diagnostic accuracy. More information on the choice of test costs is given in appendix C. Further, we evaluate the benefit of our hypothesis-driven approach in Table 3. When removing the hypothesis agent from our methodology, the decision agent has to learn to request tests and to diagnose without relying on the uncertainty-aware hypothesis generation capabilities of the hypothesis agent, explicitly trained for these objectives. The benefit of the hypothesis agent is demonstrated clearly by improvements in all metrics.
Limitations
@@
The data we are training on is retrospective with different tests missing for different patients. Furthermore, the available tests are those tests that the clinicians involved in treating that patient performed. The model can only explore a limited spread of testing pathways. It can therefore only learn to become more efficient within the testing protocols performed by doctors, e.g by skipping unnecessary test. Our objective is therefore not to discover novel clinical strategies, but to optimize decision-making within realistic, expert-demonstrated behavior. Simulation of unavailable test data could open up a pathway for modeling a more holistic clinical decision-making environment.
@@ Acknowledgements
Reproducibility Statement
In order to ensure reproducibility, we describe implementation details in Section 5 and Appendix C. Further, Appendix A provides the exact prompts used in our experiments. Lastly, we we publish our complete code containing preprocessing scripts, model code, training pipelines, and evaluation scripts at https://github.com/dharouni/LA-CDM.
Ethics Statement
The development of LA-CDM represents a step forward in AI-driven clinical decision support, offering a structured and hypothesis-driven approach to diagnosis. By iteratively requesting and interpreting clinical tests, LA-CDM has the potential to improve diagnostic accuracy while optimizing resource utilization, ultimately enhancing patient outcomes. From an ethical perspective, the deployment of LA-CDM necessitates rigorous safeguards to ensure patient safety, fairness, and transparency. As an AI system designed to support medical decision-making, it is crucial that its recommendations remain interpretable and aligned with established medical guidelines. Moreover, biases present in training data must be carefully monitored to prevent disparities in diagnostic performance across different patient populations. In the broader societal context, LA-CDM has the potential to support clinicians in environments with high cognitive load, such as emergency departments and primary care settings, where time-sensitive decision-making is critical. However, it is essential that such AI-driven tools are positioned as augmentative rather than substitutive, ensuring that they enhance rather than diminish the role of healthcare professionals. Additionally, accessibility must be prioritized to ensure equitable deployment across diverse healthcare systems, including under-resourced settings.
Appendix
@@ A.1 Hypothesis Agent Prompt
You are a helpful medical expert. Your task is to create a hypothesis for a diagnosis of a patient admitted to the emergency department complaining about abdominal symptoms. You can hypothesise one of the following four conditions:
@@
Additionally, you always have to provide your confidence as a number from 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, or 10, where 0 means very uncertain and 10 means very certain.
Always use the following format for your hypothesis:
@@
Confidence: Here, you put your confidence as a number between 0 and 10
Here are the four master rules:
1. Only answer in the above format!
2. Always provide a hypothesis, even if you are uncertain and always provide a confidence.
3. Only give one of the four possible conditions and use their exact naming.
4. Do not answer with anything other than the hypothesis and confidence as specified in the format description.
@@ 1. Only answer in the above format!
Here is the available information on the patient:
{provided_test_results}
@@ A.2 Decision Agent Prompt
To aid you in your diagnosis, you can request the results of the following tests for the patient:
@@
At the beginning of every step you will be further provided with a current hypothesis of the condition.
@@
Thought: [Here, you should think about what to do next.]
Action: [Here, you can either write "Diagnosis" if you are confident enough to provide a diagnosis or "Test" if you want to request a single test.]
Action Input: [If you chose "Diagnosis" as your action, you should put the correct condition of the four provided above here. You MUST put one of the four conditions here. If you chose "Test", provide the test name from the list mentioned above. You can only request one single test per action. Use the exact naming of the test as given above. Do not add any specification, abbreviation or explanation to the test name.]
Observation: [Here, you will receive the test results for the patient. Always write the keyword "Observation:" when requesting a test.]
The current hypothesis is: [Here you will see the current hypothesis] (confidence [Here, you will get the confidence in the hypothesis on a scale of 0 (very uncertain) to 10 (very certain)])
Thought: [Here, you evaluate the test result given above and continue with your next thought and action]
Action: [You can request more tests with the action "Test".]
Action Input: [...]
Observation: [...]
The current hypothesis is: ... (confidence: ...)
The current hypothesis is ... (confidence: ...)
Thought: [Finally, when you are confident in a diagnosis you provide your reasoning for that diagnosis first.]
Action: [You can provide a diagnosis with the action "Diagnosis"]
Action Input: [Here, you provide the condition you choose from the four given above.]
If you requested a test, the user will provide you with the test results if available. If the test is not available, do not request the same test again. If you receive an unknown test or diagnosis error, try again and ensure that you use the exact wording given above.
Here are the six master rules:
@@
2. Only ask for one of the tests given in the list above.
3. If a test returns as "not available" under no circumstance request the same test again. Try some other test. Ask for no test twice.
@@
6. Keep asking for tests until you are sufficiently confident but always provide a valid diagnosis in the end. Make sure to ask for enough tests to reach a diagnosis.
@@
{patient_history}
Appendix B Qualitative Examples
We show exemplary LA-CDM decision-making procedures. In Figure 4 and Figure 5, we show two successful predictions. In Figure 6, we show a failure case where the model gives a diagnosis prematurely without providing a reason, probably due to an overly confident hypothesis.
Figure 4: A successful diagnosis performed with the evaluation of ultrasound and CT. The model decides to confirm an initial suspicion with an ultrasound, however, the results are not conclusive, prompting another imaging test.
Figure 5: A short example diagnosis. In many cases, the model learns to correctly predict the condition very briefly with just a CT.
Figure 6: An example of a failed diagnosis of cholecystitis. The model reaches a high confidence in a wrong diagnosis too quickly and requests non-existent tests.
Appendix C Implementation Details
Dataset Pre-processing
The initial patient history is shortened by processing them with a Mixtral-8x7B (Jiang et al., 2024) model, prompted to summarize the most important aspects. Importantly, this summarization is performed without any auxiliary patient information, such as disease labels. Also the imaging reports are shortened, wherever a separation into sections was available, by only keeping the findings section of the report and removing the remaining sections. The original dataset does not have a data split, since it was intended for the evaluation of zero-shot models. We therefore split the data into a training set of 80%, and a validation and test set of 10% each.
Training Details
We use a PyTorch implementation of the Qwen-2.5-7B-Instruct model Yang et al. (2024) as base LLM and fine-tune it using LoRA Hu et al. (2022). We train our method in a cyclic training regiment with each objective trained for 100 steps in the order of (1) clinical action training, (2) hypothesis generation training, and (3) confidence calibration training. All objectives are optimized using the Adam optimizer with a learning rate of 1e51e-5 and a batch size of 2 on a single NVIDIA A40 GPU until convergence on the validation set, taking approximately 3 days.
Test Costs
Specific cost data was taken from the 2025 standard charges table sourced under the CMS HHS price transparency rule from the Beth Israel Deaconess Medical Center, where the MIMIC-CDM dataset was collected. The specific test costs are given in Table 4. For calculating the specific test punishment within the test cost reward function, all costs where normalized to sum to the reward for a correct diagnosis, such that a diagnosis pathway that requires all tests achieves a total reward of zero. Thus, it is always better to request more tests if it helps the model to reach the correct diagnosis.
Table 4: Test costs as derived from the standard charges table at the Beth Israel Deaconess Medical Center.
Test Name Cost in USD
CT 1306
MRI 4866
Radiograph 434
Ultrasound 1288
Complete Blood Count 71
Basic Metabolic Panel 298
Comprehensive Metabolic Panel 636
Renal Function Panel 394
Liver Function Panel 413
Urinalysis 50
Electrolyte Panel 134
Appendix D Experiments on Cost Sensitivity
In order to evaluate cost sensitivity, we perform two experiments: (i) Inter-institutional cost variation, and (ii) sensitivity to large price deviations.
D.1 Inter-institutional Variation
To evaluate sensitivity to inter-institutional cost variation, we create a new cost table that exhibits slight variations in pricing to our original cost table. This simulates a cost table from a different institution with different prices that, however, stay within a similar range. We provide the changed cost table in Table 5. After retraining, we observe that both diagnostic performance and test cost efficiency remain largely unaffected. The results are shown in Table 6.
D.2 Sensitivity to Specific Price Changes
In order to test model reactance to specific cost signals, we construct an artificial test table where one test’s cost, the complete blood count (CBC), has been increased to be very expensive. In this setting all requests of the CBC, lead to a test cost punishment added to the final reward of 10-10. This is significantly larger than all other test costs which are normalized to be in the range of 0 to $1$. After training with this cost table, we observe the number of CBC requests going down significantly compared to evaluating the model trained with the standard cost table. The prevalence is reduced from $19.2\%$ of cases to $2.5\%$ of cases. The diagnostic performance is also reduced as CBC can be an informative test in many situations. The F1 micro score decreases from $84.1$ to $81.9$ and the F1 macro score decreases from $81.3$ to $77.8$. This experiment clearly demonstrates the sensitivity to specific test cost distributions.
Table 5: Test costs variations are derived as slight deviations from the standard charges table at the Beth Israel Deaconess Medical Center.
Test Name Cost in USD Cost after variation in USD
CT 1306 1325
MRI 4866 4891
Radiograph 434 407
Ultrasound 1288 1274
Complete Blood Count 71 66
Basic Metabolic Panel 298 290
Comprehensive Metabolic Panel 636 626
Renal Function Panel 394 397
Liver Function Panel 413 418
Urinalysis 50 53
Electrolyte Panel 134 136
Table 6: Performance comparison of LA-CDM with and without test cost variation. (CV) marks the model with trained with the test cost variation.